OpenStudy (christina166):

Write an equation of a line perpendicular to line CD below in slope-intercept form that passes through the point (−1, 6). Point C is at (1,1) and point D is at (3,5)

OpenStudy (christina166):

@ganeshie8 Please help me!

OpenStudy (alexandervonhumboldt2):

First of all we need to find the slope of given line and its equation.

OpenStudy (christina166):

So we use the slope formula?

OpenStudy (alexandervonhumboldt2):

To do this we use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ 1-5 }{ 1-3 }=\frac{ -4 }{ -2 }=2\]

OpenStudy (alexandervonhumboldt2):

the slope of the given line is 2

OpenStudy (alexandervonhumboldt2):

slopes of perpendicular lines are negative reciprocals

OpenStudy (alexandervonhumboldt2):

what is a negative reciprocal to 2?

OpenStudy (christina166):

don't the prduct of their slopes always -1

OpenStudy (alexandervonhumboldt2):

they are but to find the equation we need to find a negative reciprocal

OpenStudy (christina166):

-0.5

OpenStudy (alexandervonhumboldt2):

yeah but we will better write it as -1/2

OpenStudy (christina166):

Oh sorry.

OpenStudy (alexandervonhumboldt2):

so we have y=mx+b y=-1/2x+b now substitute the point and solve: 6=-1/2*(-1)+b 6=1/2+b b=?

OpenStudy (christina166):

Wait what do we substitute B with, its 2 right sorry im really stupid

OpenStudy (alexandervonhumboldt2):

look: y=mx+b where m is slope: m=-1/2 y=-1/2x+b now we need to solve for y-intercept (b). to do this we substitute the point: 6=-1/2*(-1)+b b=?

OpenStudy (christina166):

OHHH I GET IT NOW THANK YOU!!!

OpenStudy (alexandervonhumboldt2):

yeah so we solve for b and then get equation y=-1/2x+?

OpenStudy (christina166):

the answer is y=2x-13

OpenStudy (alexandervonhumboldt2):

nooooooooo

OpenStudy (alexandervonhumboldt2):

why you changed -1/2 to 2?

OpenStudy (christina166):

Oh because don't we bring it over to the other side so it changes to 2

OpenStudy (alexandervonhumboldt2):

noonon we found a slope of given line: 2 then we found slope of line perp to it: -1/2 thus we have y=-1/2+b 6=-1/2*(-1)+b b=6-0.5=5.5 we get y=-1/2x+5.5

OpenStudy (christina166):

OHHHHH I feel stupid lol THANKS!!!